OpenStudy (anonymous):

What is the factored form of the expression? 4x2 – 81y2 A. (2x + 9)(2x – 9) B. (2x + 9y)(2x – 9y) C. (2x + 9y)2 D. (2x – 9y)2

OpenStudy (anonymous):

a multiple choice question ?

TheSmartOne (thesmartone):

www.mathportal.org/calculators/polynomials-solvers/polynomial-factoring-calculator.php?formId=1&val4=4x^2-81y^2

TheSmartOne (thesmartone):

Your question is in the format of (a^2-b^2)=(a+b)(a-b)

TheSmartOne (thesmartone):

In this case a would be 2x and b would be 9y So (2x+9y)(2x-9y)

OpenStudy (anonymous):

@Andresfon12 yes it is a mcq

OpenStudy (anonymous):

y= 0 X= 2

TheSmartOne (thesmartone):

@alexxkayee She needed it factored not the roots...

TheSmartOne (thesmartone):

And plus you cannot find roots because it has 2 variables..

OpenStudy (anonymous):

how would we do this then?

TheSmartOne (thesmartone):

Using this formula \[(a^2-b^2)=(a+b)(a-b) \]

TheSmartOne (thesmartone):

In this case a would be 2x and b would be 9y So \[(2x+9y)(2x-9y)=2x^2-9y^2 =4x^2-81y^2\]

TheSmartOne (thesmartone):

So basically when it is factored it is (2x+9y)(2x-9y)

OpenStudy (anonymous):

i thought it was (C) but thank you for helping me understand that its (B)

OpenStudy (anonymous):

@TheSmartOne agree, also +-=- and -+=-